OpenStudy (anonymous):

Jack is standing next to a monument in the city square. His shadow is 18 feet long. The shadow of the monument is 150 feet long. If Jack is 6 feet tall, how tall is the monument?

OpenStudy (anonymous):

6/18 = x/150, when written it is 6 over 18 and x over 150. Cross multiply. 6*150 = 18x. Divide both sides by 18. x = 50 feet. Therefore, the answer is 50 feet.